OpenStudy (anonymous):

find dy/dx if y=sin^2 x +2^(sinx)

OpenStudy (anonymous):

first part is ok right? you get \(2\sin(x)\cos(x)\) for the derivative of \(\sin^2(x)\) by the chain rule

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

as for \(2^{\sin(x)}\) this is also a chain rule problem, so long as you remember that the derivative of \(b^x\) is \(b^x\ln(b)\)

OpenStudy (anonymous):

in this case you get \[2^{\sin(x)}\ln(2)\cos(x)\]

OpenStudy (anonymous):

so thatd be 2^sinx ln (2)?

OpenStudy (anonymous):

you gotta multiply by the derivative of sine, as per the chain rule

OpenStudy (anonymous):

ok so we get 2sinxcosx + 2^sinx ln2 cos x

OpenStudy (anonymous):

yes you can factor out the cosine if you like or not

OpenStudy (anonymous):

do you factor out sinxcosx?

OpenStudy (anonymous):

your common factor is only cosine the sine in the second part is in the exponent

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok i just re-worked it out...now i see what you were talking about

OpenStudy (anonymous):

thank you very much for your help :)